OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (−2, 6) and has a slope of −1?

OpenStudy (anonymous):

the point-slope formula y2-y1=m(x2-x1)

OpenStudy (anonymous):

x + y = −4 x + y = 4 x − y = −4 x − y = 4

OpenStudy (helpblahblahblah):

B

OpenStudy (helpblahblahblah):

x+y=4

OpenStudy (anonymous):

so we have x,y and m x=-2 y=6 and m= -1

OpenStudy (anonymous):

what about On the graph of the equation 2x − y = 10, what is the value of the y-intercept?

OpenStudy (anonymous):

-10 -5 5 10

OpenStudy (anonymous):

ok well you need to know this formula too. the formula you should know the y=mx+b

OpenStudy (anonymous):

so, 2x-y=10 is given. you want it to look like y=mx+b. y=-2x+10 10 is your y-intercept

OpenStudy (anonymous):

hold!!!

OpenStudy (anonymous):

thanks not i get it. would u like a medal ?

OpenStudy (anonymous):

wait. y=2x-10

OpenStudy (anonymous):

so its -10

OpenStudy (anonymous):

yea, i hope you get it

OpenStudy (anonymous):

i do thanks